Sid the Science Kid is running a lemonade stand on a hot summer day. He wants to know how much lemonade he needs to make to sell 50 cups. He knows that each cup of lemonade requires 1/2 cup of lemon juice and 1/2 cup of sugar. He also knows that he has 2 quarts of lemon juice and 2 quarts of sugar. How many cups of lemonade can Sid make with the ingredients he has?
Answer by math_tutor2020(3817)


1 quart = 4 cups
2 quarts = 8 cups

n = number of cups of lemonade
(1/2)n = number of cups of lemon juice
(1/2)n = 8 cups
n = 8*2 = 16 cups


Answer: 16 cups of lemonade